Diffusion Morning Report 08/26/05

Ted Berg tberg at snf.stanford.edu
Fri Aug 26 08:57:15 PDT 2005


Good Morning All,
    Lots of activity today.  Nitride and LTO are back up. Poly is still 
testing.  We are drilling holes for the pumps on the new Thermco bank
Epi has leak issues. Ebeam is down for scheduled Regen.  Things are 
slowly getting back to normal.

  	Date: 	8/25/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Down 	Bank 4 move 	4.67/30
Tylan 9-Poly 	Down 	Bank 4 move 	19.05/30
Tylan 10-Nitride 	Up 	  	10.41/30
Tylan 11-BPSG-LTO 	Up 	  	5.6/30
Tylan 12-FGA 	Up 	  	 
  	  	  	 
  	  	  	 
  	  	  	 
  	  	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	Down 	Bank 4 move 	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	.
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	B on line 700 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050826/47fcf864/attachment.html>


More information about the diffusion mailing list